# Exploring the Challenge and Value of Deep Learning in Automated Skin Disease Diagnosis

### v1

@@ 5.2 Early CAD method
With the development of computer science and technologies, the CAD method was initially developed. The early development of CAD for dermatological diseases primarily focused on traditional image segmentation and feature extraction methods to analyze dermoscopic images. Initially, threshold segmentation, region growing, and edge detection techniques were widely used to identify melanoma lesions. However, these approaches relied heavily on manually extracted features, such as skin color and texture, leading to complex implementation processes, high computational costs, and poor robustness when dealing with variations in lesion appearance, uneven pigmentation, or external occlusions In the early stages, CAD systems aimed to classify dermoscopic images based on handcrafted features, improving upon traditional diagnostic methods such as the ABCD rule, Menzies method, and the seven-point checklist. Several researchers proposed advancements in this domain. Lazebnik et al. introduced a technique using approximate global geometric correspondence, segmenting input images into fine-grained subregions and computing local feature histograms, which enhanced classification accuracy [20]. Ballerini et al. developed more precise and objective feature extraction methods, reducing reliance on human-derived lesion characteristics. Their work improved the performance of the k-nearest neighbor (KNN) algorithm in dermoscopic image classification, addressing limitations such as subjectivity, inaccuracies, and weak robustness in manual feature extraction [21]. Barata et al. proposed a bag-of-features approach based on local image descriptors, using color and texture comparisons to evaluate their discriminative power in lesion classification [22]. Vasconcelos et al. enhanced CAD-based melanoma diagnosis by assessing lesion color variations according to the ABCD rule, demonstrating that color feature evaluation was highly effective for non-invasive dermoscopic assessment [23].
@@ 5.3 Deep learning method
In the field of deep learning, the data is essential, applications in medical diagnosis in particular [26]. High-resolution dermoscopic images for study could be collected from widely recognized benchmark datasets such as ISIC, HAM10000, and Derm7pt, etc. Given the challenges posed by real-world clinical settings, various preprocessing techniques are employed to enhance data quality, including image normalization, hair artifact removal, color correction, contrast enhancement, and data augmentation. These steps mitigate the impact of image noise, intra-class variability, and inter-class similarity, thereby improving model generalization.
@@
The integration of deep learning into dermatological diagnostics holds transformative potential in improving diagnostic accuracy, reducing interobserver variability, and facilitating early detection of malignant skin lesions. However, deep learning methods still face the obstacles of traditional CAD, the challenges are complex features, intra-class variation and inter-class similarity, and noise in pictures. Meanwhile, as described above, big data is critical for deep learning methods. In this situation, imbalanced classes of a dataset are a new problem to be considered.
@@ 6.3 Transfer learning
Future work should move beyond conventional fine-tuning of large-scale pre-trained models [133]. Promising directions include domain-adaptive transfer learning that incorporates dermatology-specific priors (e.g., lesion morphology and skin tone distributions) to reduce negative transfer effects, and self-supervised or few-shot transfer methods to improve performance with minimal labeled data. Cross-modal transfer learning—transferring knowledge between dermoscopic images, clinical photographs, and histopathology slides—may also improve robustness for multi-task diagnostic pipelines. Furthermore, developing lightweight transfer learning strategies optimized for deployment on edge devices would support real-time, point-of-care diagnosis. Combining transfer learning with explainable AI techniques could enhance model transparency, making it easier for clinicians to interpret transferred features and trust automated diagnostic results. These advancements are expected to further improve scalability and clinical adoption of deep learning in dermatology.
@@ 6.4 Attention mechanism
Future research should move beyond basic spatial or channel attention modules [134]. Advanced approaches such as multi-scale hierarchical attention, transformer-based self-attention, and cross-modal attention could be explored to better capture complex lesion patterns and integrate information from dermoscopic, clinical, and histopathological images. Additionally, adaptive attention mechanisms that dynamically adjust their focus based on lesion type, image quality, or patient demographics may further enhance model robustness and fairness. Combining attention with explainable AI techniques could improve interpretability and clinical trust. Exploring lightweight attention designs optimized for real-time applications on portable medical devices is another promising direction. These advancements are expected to make attention-driven models more scalable and clinically relevant in automated dermatological diagnosis.
@@ 6.6 Optimized loss function
Promising directions include dynamic loss weighting mechanisms that automatically adjust penalties based on sample difficulty, lesion rarity, or inter-class similarity during training. Combining multiple loss objectives—such as focal, Dice, contrastive, and margin-based losses—within multi-task frameworks could further improve classification and segmentation accuracy for rare dermatological conditions. In addition, integrating uncertainty-aware or noise-robust loss functions can enhance resilience to mislabeled or low-quality data, while fairness-driven loss formulations may help mitigate bias across different skin tones and demographics. Designing computationally efficient, task-specific loss functions suitable for federated and real-time applications will also be crucial for broader clinical adoption.
@@ 7 Discussion and Conclusion
The application of deep learning in dermatology has shown significant promise in improving the accuracy and efficiency of skin disease diagnosis. However, several challenges remain that must be addressed to further enhance clinical applicability and reliability. This review systematically analyzed 91 articles published between 2020 and mid-2025, identifying key obstacles including complex features, image noise, intra-class variation, inter-class similarity, and class imbalance.
@@ Acknowledgement
Support was provided by National Key Research and Development Program of China (2024YFC2511003). The authors declare that there are no conflicts of interest or competing financial interests related to this study. The authors gratefully acknowledge the constructive comments and valuable suggestions provided by the anonymous reviewers, associate editors, and the editor-in-chief, which have significantly contributed to improving the quality of this manuscript.

### v2

@@ 5.1 Traditional method
A schematic illustration of the dermoscopic diagnostic criteria is presented in Fig.7. The upper panel summarizes the ABCD rule with representative examples, and the lower panel outlines the major and minor features of the seven-point checklist. These visual elements help clarify the traditional clinical assessment methods used before deep learning–based analysis.
Fig. 7: Visual illustration of common dermoscopic diagnostic criteria.
@@ 5.2 Early CAD method
With advancements in computer science, CAD systems emerged to overcome the limitations of manual assessment. The early development of CAD for dermatological diseases primarily focused on traditional image segmentation and feature extraction methods to analyze dermoscopic images. Initially, threshold segmentation, region growing, and edge detection techniques were widely used to identify melanoma lesions. However, these approaches relied heavily on manually extracted features, such as skin color and texture, leading to complex implementation processes, high computational costs, and poor robustness when dealing with variations in lesion appearance, uneven pigmentation, or external occlusions In the early stages, CAD systems aimed to classify dermoscopic images based on handcrafted features, improving upon traditional diagnostic methods such as the ABCD rule, Menzies method, and the seven-point checklist. Several researchers proposed advancements in this domain. Lazebnik et al. introduced a technique using approximate global geometric correspondence, segmenting input images into fine-grained subregions and computing local feature histograms, which enhanced classification accuracy [20]. Ballerini et al. developed more precise and objective feature extraction methods, reducing reliance on human-derived lesion characteristics. Their work improved the performance of the k-nearest neighbor (KNN) algorithm in dermoscopic image classification, addressing limitations such as subjectivity, inaccuracies, and weak robustness in manual feature extraction [21]. Barata et al. proposed a bag-of-features approach based on local image descriptors, using color and texture comparisons to evaluate their discriminative power in lesion classification [22]. Vasconcelos et al. enhanced CAD-based melanoma diagnosis by assessing lesion color variations according to the ABCD rule, demonstrating that color feature evaluation was highly effective for non-invasive dermoscopic assessment [23].
@@ 5.3 Deep learning method
High-quality data is fundamental to deep learning, particularly in medical diagnostic applications [26]. High-resolution dermoscopic images for study could be collected from widely recognized benchmark datasets such as ISIC, HAM10000, and Derm7pt, etc. Given the challenges posed by real-world clinical settings, various preprocessing techniques are employed to enhance data quality, including image normalization, hair artifact removal, color correction, contrast enhancement, and data augmentation. These steps mitigate the impact of image noise, intra-class variability, and inter-class similarity, thereby improving model generalization.
@@
The integration of deep learning into dermatological diagnostics holds transformative potential in improving diagnostic accuracy, reducing interobserver variability, and facilitating early detection of malignant skin lesions. However, deep learning methods still encounter significant obstacles inherited from traditional CAD, including complex features, intra-class variation, and inter-class similarity. Meanwhile, as described above, big data is critical for deep learning methods. In this situation, imbalanced classes of a dataset are a new problem to be considered.
5.4 Training hyperparameters
Table 3: Summary of Training Hyperparameters.
Hyperparameter Example of Value Purpose
Optimizer Adam Stable and adaptive weight updates
Batch Size 32 Balance between memory usage and gradient stability
Learning Rate 0.001 Controls update magnitude
Scheduler StepLR Gradual LR decay for stable late-stage training
Epochs 50 Sufficient model convergence
Weight Decay 1e-5 Prevents overfitting through regularization
Training hyperparameters play a crucial role in shaping model convergence, stability, and generalization. Commonly used hyperparameters are listed in Table 3, which includes the optimizer, such as Adam and SGD, which determines how model weights are updated, and the learning rate and its scheduler, which control the step size and its decay during training. Batch size and number of epochs define how data are processed and how long the model is trained, while regularization strategies such as weight decay, dropout rates, or momentum terms help prevent overfitting and improve robustness. Additional settings, including initialization methods, gradient clipping thresholds, and early-stopping criteria, may further influence optimization behavior. Together, these hyperparameters govern the overall training dynamics and are essential for achieving stable and high-performing deep learning models.
@@ 5.6 Methods to solve image noise
Recent studies also show that noise handling should be accompanied by explicit modeling of prediction reliability. Deep models often give confident scores even when the input image contains artefacts or distribution shifts. This reduces clinical safety. Uncertainty estimation provides a way to identify low-confidence results. Common methods include Monte Carlo dropout, deep ensembles and evidential learning. These methods help separate model uncertainty and data uncertainty. Calibration is also important because raw softmax outputs usually do not reflect true confidence. Several technologies adapted this technology, such as Swin-T, Meta-UNet, Quad-Step Approach and multi-stage multi-modal learning. In the calibration model, temperature scaling and other simple post-training techniques can improve probability calibration. Reliability diagrams and Brier scores are often used to assess calibration quality. Integrating uncertainty estimation with noise reduction can improve the trustworthiness of skin lesion analysis and support safer clinical use.
@@ 6.3 Transfer learning
Future work should move beyond conventional fine-tuning of large-scale pre-trained models [133]. Promising directions include domain-adaptive transfer learning that incorporates dermatology-specific priors to reduce negative transfer effects, and self-supervised or few-shot transfer methods to improve performance with minimal labeled data. Cross-modal transfer learning—transferring knowledge between dermoscopic images, clinical photographs, and histopathology slides—may also improve robustness for multi-task diagnostic pipelines. Furthermore, developing lightweight transfer learning strategies optimized for deployment on edge devices would support real-time, point-of-care diagnosis. Combining transfer learning with explainable AI techniques could enhance model transparency, making it easier for clinicians to interpret transferred features and trust automated diagnostic results. These advancements are expected to further improve scalability and clinical adoption of deep learning in dermatology.
Moreover, when applying transfer learning to new datasets or imaging conditions, a key challenge is to ensure cross-dataset generalization under domain shift to maintain robust performance across varying data distributions. In recent years, the challenge of cross-dataset generalization and domain shift in dermatological image analysis has become increasingly prominent. For example, a new dataset called IMPS has been created by combining diverse clinical image sources via a new dataset. After applying Domain‑Adversarial Neural Network, an unsupervised domain adaption method, the generalisation could be enhanced [134]. In addition, the DRMv2Net model also demonstrates strong cross-dataset generalization. This model is evaluated on both the ISIC 2357 and PAD-UFES 20 datasets and achieving accuracies of 96.11% and 96.17% respectively, thereby indicating robustness against domain shift across different imaging sources. [135]. Prompt-based methods built on global and local prompts can effectively separate domain-shared and domain-specific knowledge. They achieve superior segmentation generalization across five heterogeneous datasets [136]. Moreover, Fourier-domain style transfer [137] and structure-preserving image translation [138] further enhance model adaptability and robustness under varying imaging protocols, devices, and magnification levels [139].
In a word, the convergence of transfer learning and cross-dataset generalization presents a promising frontier for dermatological AI. Transfer learning enables the reuse of pretrained representations and accelerates model development with limited labeled data, while cross-dataset generalization addresses the critical problem of domain shift when models encounter new imaging sources or populations.
@@ 6.4 Attention mechanism
Future research should move beyond basic spatial or channel attention modules [140, 141]. Advanced approaches such as multi-scale hierarchical attention, transformer-based self-attention, and cross-modal attention could be explored to better capture complex lesion patterns and integrate information from dermoscopic, clinical, and histopathological images. Additionally, adaptive attention mechanisms that dynamically adjust their focus based on lesion type, image quality, or patient demographics may further enhance model robustness and fairness. Combining attention with explainable AI techniques could improve interpretability and clinical trust. Exploring lightweight attention designs optimized for real-time applications on portable medical devices is another promising direction. These advancements are expected to make attention-driven models more scalable and clinically relevant in automated dermatological diagnosis.
In addition to the studies originally reviewed, we further examined recent advances that explicitly integrate hierarchical attention with multi-scale feature design, which aligns closely with the reviewers’ suggestions. For example, AMST-Net introduces a dual-encoder architecture in which a multi-scale Transformer generates hierarchical attention features that are progressively fused with CNN representations, demonstrating the effectiveness of combining layered self-attention with multi-resolution information for improving boundary precision and global context modeling. These complementary findings reinforce the growing trend toward adaptive, multi-level attention frameworks in dermatological imaging and support our expanded discussion on attention hierarchies and multi-scale designs in automated skin lesion analysis [142]. The MTA-Net model explicitly integrates multi-scale triplet spatial and channel attention, enabling the network to capture fine-grained lesion characteristics across different scales and dimensions. By enriching both spatial and channel-wise representations, MTA-Net significantly improves multiclass skin lesion classification performance on HAM10000 and ISIC2019, demonstrating the effectiveness of hierarchical multi-scale attention in handling high inter-class similarity and intra-class variability [143]. HMSA-Net introduces a hierarchical multi-scale aggregation framework that integrates Res2Net-based multi-scale contextual encoding with Hierarchical Attention Refinement Blocks (HARBs), enabling dynamic channel recalibration through squeeze-and-excitation mechanisms. By fusing features across multiple resolutions and emphasizing anatomically critical structures, the model demonstrates strong segmentation performance across diverse biomedical imaging modalities, including ISIC-2018 skin lesion segmentation [144]. There are also several studies proves that multi-scale attention mechanisms, including structures such as multi-scale spatial pyramid attention [145, 146] and hierarchical attention modules [147], effectively fuse features across scales and levels. They enhance the model’s ability to recognize complex skin lesions. By capturing both local details and global context, these mechanisms improve detection of small lesions and blurred boundaries, and they strengthen the model’s generalization across diverse disease types.
Overall, these recent developments illustrate that combining hierarchical attention with multi-scale feature aggregation provides a powerful mechanism for capturing both global contextual cues and subtle local lesion details. This convergence of attention-driven and multi-scale design principles is becoming a key direction for building more robust, generalizable, and clinically reliable dermatological deep learning models.
@@ 6.6 Optimized loss function
Promising directions include dynamic loss weighting mechanisms that automatically adjust penalties based on sample difficulty, lesion rarity, or inter-class similarity during training. Combining multiple loss objectives—such as focal, Dice, contrastive, and margin-based losses—within multi-task frameworks could further improve classification and segmentation accuracy for rare dermatological conditions. For instance, dice-based loss functions are also widely adopted to mitigate class imbalance, especially in lesion segmentation tasks where the foreground region is small relative to the background. In addition, integrating uncertainty-aware or noise-robust loss functions can enhance resilience to mislabeled or low-quality data, while fairness-driven loss formulations may help mitigate bias across different skin tones and demographics. Designing computationally efficient, task-specific loss functions suitable for federated and real-time applications will also be crucial for broader clinical adoption.
7 Frontier Advances and Emerging Paradigms
The rapid evolution of dermatological artificial intelligence in recent years has given rise to a new generation of frontier technologies that extend far beyond conventional deep learning solutions. These emerging paradigms integrate multi-dataset fusion, cross-domain generalization, attention-driven hybrid architectures, and uncertainty-aware learning, collectively reshaping the landscape of automated skin disease diagnosis. Unlike earlier methods that rely on single-source data or isolated network designs, current advances emphasize robustness, adaptability, and clinical reliability through the integration of diverse feature sources, hierarchical attention mechanisms, and domain-shift-resistant modeling strategies. This section highlights these transformative developments, outlining how they mark a paradigm shift toward more comprehensive, scalable, and clinically deployable diagnostic systems.
7.1 Multimodal and Cross-Domain Fusion
Recent work has shown that multi-dataset fusion and cross-domain learning can substantially enhance the robustness of dermatological deep learning models. A representative example is EffiCAT (EfficientNet Concatenation Attention Technology), a hybrid architecture designed to improve generalization across heterogeneous datasets. EffiCAT integrates complementary feature representations extracted from EfficientNet-B0 and EfficientNet-B4 through feature concatenation, followed by multiple attention modules including a dual channel attention layer and CBAM to refine salient lesion characteristics. EffiCAT demonstrates strong performance, achieving 94.48% accuracy on the merged dataset and maintaining high cross-dataset robustness with 92.08% accuracy on ISIC-2018 [149]. Recent studies beyond EffiCAT have further demonstrated the value of multimodal and cross-domain fusion in dermatological AI. Several works jointly model dermoscopic images, clinical photographs, and patient metadata through attention-based or multi-stage fusion strategies, such as adversarial attention fusion networks, tri-modal cross-attention transformers (TMCT), and hierarchical fusion architectures like MsfuseNet [150]. Cai et al. also built a hybrid skin cancer segmentation model that integrates CNNs, Transformers, and a U-Net framework with a dual-frequency cross-fusion strategy.By combining features extracted from the Discrete Walsh Transform and the Haar wavelet transform, the model leverages both high-frequency details and low-frequency structural information, while an attention mechanism enhances the fusion of these complementary features [151]. Recent studies also show that fusing skin-disease images with patient metadata can markedly improve accuracy and robustness. In specific, cross-attention mechanisms [152], mutual-information switching modules [153], transformer architectures [154], enable dynamic integration of heterogeneous features and capture richer lesion information. Some models now achieve over 95% accuracy on public datasets such as HAM10000 and ISIC 2018, with reports approaching 99
Multimodal and cross-domain fusion carries substantial technical importance in advancing dermatological artificial intelligence [155]. By integrating heterogeneous information sources such as dermoscopic images, clinical photographs, and patient-level metadata, fusion-based models capture complementary feature representations that cannot be derived from any single modality alone. This integrated learning process improves resilience to domain shifts. Meanwhile, dataset-specific bias, and enhances generalization across diverse clinical environments can also be reduced by this technology. Moreover, cross-domain alignment enables models to learn invariant patterns shared across imaging conditions, acquisition devices, and patient populations, thereby strengthening diagnostic reliability.
7.2 Uncertainty-aware and Reliability-calibrated Learning
Although interpretability has gained increasing attention, recent studies emphasize that reliability assessment. For instance, uncertainty estimation in deep-learning models solving skin lesion recognition enables the prediction not only of labels but also of the model’s confidence, distinguishing between data-driven uncertainty and model-driven uncertainty [156],[157].
Beyond estimating uncertainty, it is vital to ensure that confidence scores are calibrated. For example, in A novel uncertainty‑aware deep learning technique with an application on skin cancer diagnosis the authors develop a Monte Carlo Dropout (MCD) based framework augmented with an uncertainty-aware loss which seeks to reduce expected calibration error (ECE) and improve uncertainty accuracy (UA) [158]. These findings underscore that for clinical deployment, systems must not only achieve high accuracy, but also provide reliable, calibrated confidence estimates to support risk-aware decision making. By enabling reliable confidence assessment, uncertainty-aware models help identify cases where predictions may be untrustworthy and thus trigger expert review, which is essential for safe clinical deployment of AI systems.
Besides, the combination of multimodal fusion and diagnostic uncertainty quantification further improves accuracy, reliability, and clinical utility. This integration provides a solid foundation for the safe deployment of intelligent dermatological diagnosis [159, 160].
@@ 8 Discussion and Conclusion
8.1 Historical progression
The Fig.11 illustrates the methodological evolution from rule-based clinical assessment to early computer-aided diagnosis using handcrafted features, and finally to modern deep learning and explainable AI models that integrate attention mechanisms, multimodal fusion, and interpretability tools such as Grad-CAM and SHAP.
This historical progression highlights that performance gains alone are not sufficient for clinically reliable skin disease diagnosis. Traditional rule-based methods offered transparency but lacked robustness, whereas early CAD systems and deep learning models improved accuracy at the cost of increased opacity. Recent advances in explainable and uncertainty-aware AI suggest that future work must balance accuracy with interpretability, calibration, and cross-domain robustness so that automated systems can truly support dermatologists in safe, trustworthy clinical decision-making.
Fig. 11: Progression from traditional diagnostic techniques to deep learning.
8.2 Technological summary
Now, the application of deep learning in dermatology has shown significant promise in improving the accuracy and efficiency of skin disease diagnosis. However, several challenges remain that must be addressed to further enhance clinical applicability and reliability. This review systematically analyzed 91 articles published between 2020 and mid-2025, identifying key obstacles including complex features, image noise, intra-class variation, inter-class similarity, and class imbalance.
Table 8: Distribution of studies across four key challenges in automated skin disease diagnosis.
Research Theme Number of Studies
Complex Features 18
Image Noise 21
Intra-Class Variation and Inter-Class Similarity 23
Class Imbalance 29
The Table 8 presents the distribution of 91 studies across four major challenges in automated skin disease diagnosis: complex features, image noise, intra-class variation, and inter-class similarity. Notably, the largest number of studies (29) addresses class imbalance, indicating that this issue is a primary concern in the field. A significant portion of the studies (23) also focuses on intra-class variation and inter-class similarity, highlighting the difficulty in distinguishing lesions within the same class. Image noise is another common challenge, with 21 studies dedicated to mitigating it. In contrast, complex features received comparatively less attention, with only 18 studies focusing on this area, suggesting that it remains a more complex and perhaps less explored issue in dermatological deep learning applications.
@@ 8.2 Technological summary
8.3 Comments
Methods addressing complex features benefit from hybrid CNN-Transformer architectures, multi-scale feature fusion, and attention-based representations, which collectively enhance the extraction of irregular lesion shapes, color variations, and global contextual cues. These models achieve strong performance but often require high computational resources and large datasets. Their interpretability mainly relies on attention maps or CAM-based visualizations, which provide partial but not fully transparent insights into decision-making.
Image noise is targeted by technologies including preprocessing pipelines, contrast enhancement, denoising filters, and knowledge-distillation-based light-weighted models-effectively improve robustness to artifacts such as hair, illumination variations, and background clutter. However, they frequently depend on handcrafted preprocessing steps, which may limit generalizability across devices or clinical environments. Interpretability remains limited because many enhancements occur at the pixel or feature level, which is less intuitive for clinicians.
Approaches for intra-class variation and inter-class similarity utilize attention mechanisms, multi-scale learning, metadata fusion, and hierarchical classification strategies to better capture subtle local structures and reduce ambiguity between visually similar lesion types. These designs are clinically relevant, as they mirror the multi-scale visual reasoning used by dermatologists. However, the increased architectural complexity may reduce model transparency. Attention-based explanations enhance interpretability, but they do not necessarily ensure causal clinical correctness.
Methods addressing class imbalance rely on data augmentation, GAN or diffusion-based synthesis, weighted loss functions, and resampling strategies. Such techniques substantially improve minority-class sensitivity and prevent models from being biased toward common lesion types. Yet, synthetic samples may introduce distributional shifts or unrealistic artifacts, and weighted losses can over-correct the imbalance. Interpretability is generally not directly affected, but clinical reliability requires verification to ensure that augmented images do not mislead the model.
Overall, as listed in Table 9, recent advances show strong technical performance, their practical and clinical reliability depends on consistent interpretability, robustness across diverse patient groups, and validation under real-world conditions. These considerations remain essential for translating deep learning systems into trustworthy dermatological decision support tools.
Table 9: Compact comparative matrix of deep-learning methods addressing 4 major challenges.
Challenge Representative Techniques Performance Robustness Interpretability Clinical Reliability
Complex Feature Representation CNN–Transformer hybrids; multi-scale fusion; attention mechanisms Strong: typically higher accuracy than CNN baselines Moderate Weak Moderate
Noise & Artifact Robustness Preprocessing pipelines; contrast enhancement; denoising; knowledge distillation Moderate: mainly improves performance under noisy inputs Strong Weak Weak
Fine-grained Visual Discrimination (Intra-/Inter-class variation) Multi-scale learning; hierarchical attention; metadata fusion Strong: better separation of visually similar lesions Moderate Weak Moderate
Class Imbalance Data augmentation; GAN/diffusion synthesis; weighted loss; resampling Moderate: higher minority-class sensitivity; overall metrics vary Moderate Moderate Weak
8.4 Conclusion
Overall, this study makes a distinctive contribution by establishing a comprehensive, PRISMA-based challenge–solution taxonomy for deep learning in dermatology, bridging technical innovation with clinical applicability. This framework may serve as a foundation for future research toward reliable, interpretable, and equitable dermatological AI systems.
@@ Acknowledgement
Support was provided by National Key Research and Development Program of China[http://dx.doi.org/10.13039/501100013290] (Grant No. 2024YFC2511003) and Natural Science Foundation of Zhejiang Province (Project No. LDT23E05015A01). The authors declare that there are no conflicts of interest or competing financial interests related to this study. The authors gratefully acknowledge the constructive comments and valuable suggestions provided by the anonymous reviewers, associate editors, and the editor-in-chief, which have significantly contributed to improving the quality of this manuscript.
This work was previously available as a preprint on arXiv.
https://arxiv.org/abs/2510.03869
Appendix
Checklist Item Location in Paper
Title
Identify the report as a systematic review. The title clearly states that it is a systematic review of deep learning in automated skin disease diagnosis.
Abstract
See the PRISMA 2020 for Abstracts checklist. The abstract includes background, objective, methods, results, and conclusion.
Introduction
Describe the rationale for the review in the context of existing knowledge. The introduction describes the challenges in automated skin disease diagnosis and why this review is needed.
Provide an explicit statement of the objective(s) or question(s) the review addresses. The objective of the review is stated as evaluating deep learning approaches for automated skin disease diagnosis.
Methods
Specify the inclusion and exclusion criteria for the review. The eligibility criteria for included studies are detailed, including the year range (2020–2025) and the focus on deep learning methods for skin disease diagnosis.
Specify all databases, registers, websites, organisations, reference lists, and other sources searched or consulted to identify studies. Databases such as PubMed, IEEE Xplore, and Web of Science are listed with the date of last search.
Present the full search strategies for all databases. The full search strategies for the included databases are detailed.
Specify the methods used to decide whether a study met the inclusion criteria. The study selection process is explained, with a PRISMA flow diagram included in Section 2.
Specify the methods used to collect data from reports. Data extraction methods are outlined, including details on how data were reviewed and confirmed by multiple reviewers in Section 2.
List and define all outcomes for which data were sought. The outcomes such as diagnostic accuracy (e.g., sensitivity, specificity) are described in Section 2.
Specify the methods used to assess risk of bias. Risk of bias assessment using tools like the Newcastle-Ottawa Scale is described in Section 2.
Specify for each outcome the effect measure(s) used. Effect measures like accuracy, precision, and recall are specified for each outcome in Section 2.
Describe the processes used to decide which studies were eligible for each synthesis. The synthesis methodology is explained in Section 2, describing how studies were grouped and compared.
Describe any methods used to explore possible causes of heterogeneity among study results. Methods like subgroup analysis and meta-regression are discussed in Section 2.
Describe any sensitivity analyses conducted to assess robustness of the synthesized results. Given the methodological homogeneity of the included studies and their similar approaches to deep learning for automated skin disease diagnosis, a sensitivity analysis was not deemed necessary. The results were consistent across studies, and no significant variation warranted further testing.
Results
Describe the results of the search and selection process, from the number of records identified in the search to the number of studies included in the review. A flowchart is provided detailing the number of records screened and studies included, as shown in Fig. 2 in Section 3.
Cite studies that might appear to meet the inclusion criteria, but which were excluded, and explain why they were excluded. The reasons for excluding studies are described in Section 2.
Cite each included study and present its characteristics. Study characteristics are summarized in a table, detailing each study’s design, deep learning methods, and outcomes in Section 5.
Present assessments of risk of bias for each included study. Risk of bias for each study is assessed and presented in the results section, Section 2.
For all outcomes, present summary statistics for each group and effect estimate and its precision. We did not calculate summary statistics or effect estimates for the outcomes. Instead, the findings are described thematically based on the individual studies included.
For each synthesis, briefly summarise the characteristics and risk of bias among contributing studies. Study characteristics and risk of bias are summarized for each synthesis in Section 5.
Present results of all statistical syntheses conducted. Because this review synthesizes the literature qualitatively rather than through statistical aggregation, no formal statistical synthesis or meta-analytic summary estimates were produced. The findings are integrated descriptively, without calculating pooled effects, confidence intervals, or measures of heterogeneity.
Present results of all investigations of possible causes of heterogeneity among study results. Heterogeneity analysis results are presented in Section 8.
Present results of all sensitivity analyses conducted to assess the robustness of the synthesized results. No sensitivity analyses were performed in this review; therefore, no corresponding results are reported.
Discussion
Provide a general interpretation of the results in the context of other evidence. The discussion section interprets the findings in the context of existing literature in Section 8.
Discuss any limitations of the evidence included in the review. The limitations of the included studies and the review process are discussed in Section 8.
Discuss any limitations of the review processes used. The limitations in the review methodology are explained in Section 8.
Discuss implications of the results for practice, policy, and future research. The practical implications for clinical practice and future research directions are discussed in Section 8.